OpenStudy (anonymous):

point/slope form. Write an equation in point slope form y-y = m(x-x) with point (-2,5) m = -6 ... Then put in Standard Form Ax+By=C

OpenStudy (ajprincess):

did u try this and vat r u gettng?

OpenStudy (anonymous):

I'm confused as to how to figure out the problem. I get as far as putting it in point/slope form of .... y-5=-6(x-(-2)) I just don't know how to figure problem to put in Standard Form of Ax+By=C

OpenStudy (ajprincess):

Nw remove the brackets and simplify

OpenStudy (anonymous):

y-5 = -6x+2

OpenStudy (anonymous):

wait sorry .... its y-5=-6x-12

OpenStudy (ajprincess):

Nw bring the terms with x and y to one side and the numbers to the other side.

OpenStudy (anonymous):

ok so we subtract y from each side and add 12 to each side which then gives me 7=-6x-y

OpenStudy (anonymous):

the reason i'm stuck is because we cannot have a -x in standard form so I'm not sure how to make into a positive

OpenStudy (ajprincess):

multiply by negative both the sides.

OpenStudy (anonymous):

would we just divide the -6x by -1 and do the same to the 7 making it -7?

OpenStudy (anonymous):

the y is left alone ... then answer is 6x-y=-7 Yes?

OpenStudy (ajprincess):

u shud divide y also by -1

OpenStudy (anonymous):

we do? so it's 6x+y=-7 ? Why do we divide y also?

OpenStudy (ajprincess):

when u r dividing by a number both sides u need to divide all the terms nt just one. do u get it?

OpenStudy (anonymous):

yes thank u

OpenStudy (ajprincess):

welcome.:)